OpenStudy (anonymous):

PLEASE HELP ME!!!!!? Determine which system below will produce infinitely many solutions. -6x + 3y = 18 4x - 3y = 6 2x + 4y = 24 6x + 12y = 36 3x - y = 14 -9x + 3y = -42 5x + 2y = 13 -x + 4y = -6

OpenStudy (godlovesme):

An equation have infinitely many solution when it have the same slope and y-intercept can u change the equations to slope-intercept form? And WELCOME TO OS :D

OpenStudy (anonymous):

ummmm....... I'm not good at doing that. @Godlovesme

OpenStudy (godlovesme):

I wish I could write it all for u but I have gtg I'll tag someone that could help u sorry @jim-thompso

OpenStudy (anonymous):

thank you! @Godlovesme

OpenStudy (anonymous):

hi everyone I am new here I need help

OpenStudy (godlovesme):

NP :)I really really apologize =0

OpenStudy (anonymous):

can you help me @Preetha ? Its ok if I need helpm with something else in the future ill ask @Godlovesme

OpenStudy (godlovesme):

K I would suggest u to use geogebra which helps u graph and u canc whenu enter the eqns and if oonly one line pops up...BINGO! LOLL xD thanks for the medal too!!

OpenStudy (anonymous):

ok thanks! @Godlovesme

jimthompson5910 (jim_thompson5910):

Godlovesme has a good suggestion. If you graph a system and you get 1 line showing up, which in reality is a situation where 1 line overlaps another perfectly (covering it up), then you'll have a dependent system with infinitely many solutions.

OpenStudy (anonymous):

its not letting me down load it for some reason! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's ok, there's always desmos https://www.desmos.com/calculator it's an online calculator you don't need to download

jimthompson5910 (jim_thompson5910):

graph each system one at a time

OpenStudy (anonymous):

ok so i know a isnt it. Am i correct so far? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes choice A should have 2 lines that cross at exactly 1 point

jimthompson5910 (jim_thompson5910):

that crossing point, or intersection point, is the only solution so there aren't infinitely many solutions for choice A

OpenStudy (anonymous):

wait is is c then @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you are correct

jimthompson5910 (jim_thompson5910):

that system should produce 1 line (2 lines really, one hidden below another)

OpenStudy (anonymous):

because there is infinitely many solutions for it because both lines are exactly the same!

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

thank you for sending that link! that helped a lot!!!!!! :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm glad it did

OpenStudy (anonymous):

do you think you can help with one more!? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 a.) x=0 b.) x=2 c.) there is no x value, as there is no solution d.) x can be any value, as there is infinitely many solutions @jim_thompson5910

jimthompson5910 (jim_thompson5910):

are you able to solve 2x + y = 1 for y? ie are you able to isolate y in the equation 2x + y = 1 ?

OpenStudy (anonymous):

yes you can.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you get when you solve for y

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

subtract 2x from both sides 2x + y = 1 2x + y-2x = 1 -2x y = -2x + 1

jimthompson5910 (jim_thompson5910):

hopefully you see how that works?

OpenStudy (anonymous):

Yes i did it differently i forgot that way

jimthompson5910 (jim_thompson5910):

now you focus on the second equation and replace y with -2x+1 4x + 2y = -1 4x + 2(-2x+1) = -1 ... replace y with -2x+1 solve for x

OpenStudy (anonymous):

ok so there is no solution. am i correct? therefore the answer is c.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct, you should have 2 = -1 when you simplify

OpenStudy (anonymous):

yep thats what i got!!!!!!!!!!!!!!! thank you sooooo much you made this clear to me

jimthompson5910 (jim_thompson5910):

you're welcome